OpenStudy (owl2023):

Confused on this. Euler's Formula I know E=24 and V=15 I got 11 for number of faces? Yet people are saying it's 9... please help

OpenStudy (owl2023):

@MysticFury

OpenStudy (mysticfury):

This page lists proofs of the Euler formula: for any convex polyhedron, the number of vertices and faces together is exactly two more than the number of edges. Symbolically V−E+F=2.

OpenStudy (owl2023):

Yes I did that and got 11... Yet a lot of my friends who did the homework before me are saying the correct answer is 9? I'm really confused if they are or wrong or I'm wrong.

OpenStudy (mysticfury):

thats what i found for you, hope it helps a little, sorry im not sure on this :/ i havent learned this yet lol

OpenStudy (owl2023):

That's okay, Do you happen to know anyone that may be able to help me? Could you tag them or try to get them here. I really thank you for the effort.

OpenStudy (mysticfury):

@YoungStudier @Qwertty123 @Aveline @mathstudent55 @rebeccaxhawaii @ILovePuppiesLol @Awolflover1 @gabbyalicorn one of them has to be able to help :) and np i tried :P

OpenStudy (owl2023):

I thought EULER'S FORMULA was just F+V=E+2 Yet I get 11? People say it's 9... I don't know what to do?

OpenStudy (aveline):

Sorry, not familiar with this concept >_<